# Beyond algorithm hyperparameters: on preprocessing hyperparameters and associated pitfalls in machine learning applications

### v1

@@ 4.1.3 Tuning procedures
When tuning an HP $\lambda_{j}$, it is often not reasonable to consider all possible HP values (i.e. all values in $\Lambda_{j}$). For example, this applies if certain values of $\lambda_{j}$ are already known to cause overfitting or convergence issues. Moreover, when $\lambda_{j}$ is a preprocessing HP, $\Lambda_{j}$ may not even be formally specified (see Section 2.3.3). To perform HP tuning, it is thus essential to specify a search space $\tilde{\Lambda}_{j}$ for each HP, where $\tilde{\Lambda}_{j}$ is a bounded subset of $\Lambda_{j}$ and determines the HP values that are considered for tuning (Bischl et al.,, 2023). For example, if the HPs of the CART algorithm, $\lambda_{cp}$ and $\lambda_{minsplit}$ with $\Lambda_{cp}=[0,1]$ and $\Lambda_{minbucket}=\{1,\dots,n_{\text{train}}\}$, are tuned, their search spaces could be defined as $\tilde{\Lambda}_{cp}=[0.001,0.1]$ and $\tilde{\Lambda}_{minbucket}=\{5,\dots,25\}$. The (overall) search space of all $J$ HPs is denoted as $\boldsymbol{\tilde{\Lambda}}=\tilde{\Lambda}_{1}\times\dots\times\tilde{\Lambda}_{J}$.
@@
Choosing a search space involves the following trade-off: If the search space is too small, the prediction error achieved by $\boldsymbol{\tilde{\lambda}}^{\ast}$ and $\boldsymbol{\lambda}^{\ast}$ may differ greatly. On the other hand, if the search space is too large, this decreases the chance of finding $\boldsymbol{\tilde{\lambda}}^{\ast}$ (or a HP configuration that leads to a comparable prediction error) within a given computational budget (Bischl et al.,, 2023).
@@
Unless the specified search space $\boldsymbol{\tilde{\Lambda}}$ is very small such as when only a few categorical HPs are tuned, evaluating all HP configurations in the search space can be computationally challenging, or even infeasible. For example, even if $\lambda_{cp}$ and $\lambda_{minbucket}$ are the only HPs being tuned, with the search spaces as specified above and $\tilde{\Lambda}_{cp}$ being searched in increments of 0.001, $C=100\times 21=2100$ candidate HP configurations would need to be evaluated. Accordingly, one or several criteria must be specified to terminate the tuning process once it is met. The trade-off to consider when choosing a termination criterion is that the tuning process should neither stop before finding $\boldsymbol{\tilde{\lambda}}^{\ast}$, nor should it continue longer than necessary, which would result in an inefficient use of resources and, as we will discuss below, increase the risk of overtuning (Bischl et al.,, 2023).
In automated tuning procedures, commonly-used criteria are based on the number of evaluations or the runtime. However, additional criteria such as reaching a certain performance level or stagnation of performance might also be reasonable (Bartz et al.,, 2023; Bischl et al.,, 2023). Similar termination criteria, though often more intuitive than formally specified, may also exist for manual tuning when, for example, the user stops searching when satisfied by the reached performance level or gives up searching after a certain amount of time.
@@
Since in many cases, only a subset of all HP configurations in the search space can be evaluated before the tuning process is terminated, the way in which the sequence of evaluations is determined, also called search strategy or HPO algorithm (Elsken et al.,, 2019; Bischl et al.,, 2023), is another important component of the tuning procedure. Search strategies can be characterized by several aspects, such as the amount of time they spend inferring new candidate HP configurations from already evaluated ones (known as the inference vs. search trade-off; Bischl et al.,, 2023). For example, search strategies such as evolutionary algorithms and Bayesian optimization consider the distribution and results of previously evaluated HP configurations to propose new configurations. In contrast, the commonly used random search strategy simply draws HP configurations from a predefined, typically uniform, distribution without taking into account past evaluations (see, e.g., Feurer and Hutter,, 2019, Bischl et al.,, 2023, or Bartz et al.,, 2023 for more details and other search strategies). In the special case where only the set of selected features is tuned, a well-known automated search strategy is backward or forward feature selection (see, e.g., Hastie et al.,, 2009).
@@
In automated tuning procedures, all HPs are usually tuned jointly, i.e. each evaluated HP configuration potentially considers different values of each HP. However, the HPs could also be tuned sequentially, i.e. the complete tuning procedure is repeated for each HP (Probst et al.,, 2019; Waldron et al.,, 2011). For example, in a setting with three HPs (i.e. $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\lambda_{3})$), $\lambda_{1}$ would be tuned first with $\lambda_{2}$ and $\lambda_{3}$ set to default, which yields $\lambda_{1}^{\mathrm{II}}$. Then, $\lambda_{2}$ is tuned with $\lambda_{1}=\lambda_{1}^{\mathrm{II}}$ and $\lambda_{3}$ set to its default. Finally, $\lambda_{3}$ is tuned with $\lambda_{1}=\lambda_{1}^{\mathrm{II}}$ and $\lambda_{2}=\lambda_{2}^{\mathrm{II}}$, yielding $\lambda_{3}^{\mathrm{II}}$. As sequential tuning does not consider any interaction effects between the HPs, it is generally less likely to yield a $\boldsymbol{\lambda}^{\mathrm{II}}$ comparable to $\boldsymbol{\tilde{\lambda}}^{\ast}$ than joint tuning. On the other hand, sequential tuning demands less time, with the maximum number of evaluations increasing linearly rather than exponentially with the number of HPs to tune, as is the case with joint tuning. Hence, it could be a realistic approach for manual tuning.
@@
Due to the optimistic bias of the apparent error, the standard approach for automated HP tuning is to employ a resampling method. In the case of $k$-fold CV, which is a common choice for HP tuning (Bischl et al.,, 2023), this means that for each candidate HP configuration $\boldsymbol{\lambda}^{(c)}$, $k$ models are trained and evaluated on different subsets of $\mathcal{D}_{\text{train}}$.
@@
While resampling methods provide an improvement over using the apparent error, the corresponding estimators also exhibit a certain degree of pessimistic bias and variance (with the degree of bias and variance depending on the resampling method used, as discussed in Section 3.2.1). A potential pitfall arising from the variance is that the winning HP configuration, $\boldsymbol{\lambda}^{\mathrm{II}}$, may have been selected simply because the trained prediction model(s) using $\boldsymbol{\lambda}^{\mathrm{II}}$ performed particularly well by chance on the specified test data set(s) $\mathcal{D}_{\text{test}}$, which are the same for each evaluated HP configuration. This means that the HP selection has essentially been overfitted to the respective test data set(s) $\mathcal{D}_{\text{test}}$, which in this context is also referred to as overtuning, overhyping, or oversearching (Hosseini et al.,, 2020; Quinlan and Cameron-Jones,, 1995; Ng,, 1997; Feurer and Hutter,, 2019; Bischl et al.,, 2023; Cawley and Talbot,, 2010). If the true prediction error of $\boldsymbol{\lambda}^{\mathrm{II}}$ is still comparable to the prediction error of $\boldsymbol{\tilde{\lambda}}^{\ast}$, overtuning effects are negligible. However, there might also be scenarios in which the true prediction error of $\boldsymbol{\lambda}^{\mathrm{II}}$ is no better, or even worse, than that of the default HP configuration, but its estimated prediction error is drastically deflated (i.e. over-optimistic), as the corresponding prediction model(s) that were trained during resampling incidentally fit very well to the specific noise pattern in the respective test data set(s) $\mathcal{D}_{\text{test}}$. This has been demonstrated in several experiments where tuning was conducted on null data (i.e. data without any true signal), yet the prediction error estimate of the selected HP configuration $\boldsymbol{\lambda}^{\mathrm{II}}$ was substantially smaller (i.e. better) than its true prediction error indicating random prediction (Hosseini et al.,, 2020; Varma and Simon,, 2006; Boulesteix and Strobl,, 2009; Bischl et al.,, 2023).
@@
Note that since the HPs are overfitted to the test data set(s) $\mathcal{D}_{\text{test}}$, which are not seen during training on the corresponding $\mathcal{D}^{\prime}_{\text{train}}$, overtuning occurs on a higher level than overfitting of the model parameters (see Section 3.1). Accordingly, overtuning effects may only be visible after evaluating a large number of HP configurations (Bischl et al.,, 2023). However, literature suggests that the risk of overtuning does not only depend on the number of evaluated HP configurations but also for example on the search strategy, the type of tuned HP, and the number of observations in $\mathcal{D}_{\text{train}}$ (Hosseini et al.,, 2020; Wainer and Cawley,, 2021; Cawley and Talbot,, 2010). In general, overtuning is considered an open problem of HP tuning, and although strategies have been suggested to avoid it (e.g., using different splits for each evaluation, Nagler et al.,, 2024), there are no commonly agreed-upon solutions (Feurer and Hutter,, 2019).
@@ 4.2.1 Evaluation of a model generated on all available data
As in Setting I, reporting the apparent error for model evaluation is inappropriate in Setting II (see Figure 3, model evaluation a). In this case, however, the designated test data set $\mathcal{D}_{\text{test}}=\mathcal{D}_{\text{train}}=\mathcal{D}$ is even used twice during model generation: first during the HP tuning process and then again in the final training process. Depending on the specific tuning procedure employed, this can introduce an even greater optimistic bias compared to, for example, using default HP values. Although the apparent error is generally not suitable for assessing a model’s performance, some users who performed tuning via resampling may mistakenly believe it now reflects a form of resampling error. This was noted by Neunhoeffer and Sternberg, (2019), who also reference a paper that appears to have fallen into this pitfall.
@@
Similar to Setting I, an alternative evaluation procedure in Setting II is to employ a resampling method (see Figure 3, model evaluation b). In principle, the chosen resampling method is carried out as described in Section 3.2.1, except that in each resampling iteration, the model is trained on $\mathcal{D}^{\prime}_{\text{train}}$ and evaluated on $\mathcal{D}_{\text{test}}$ with $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{II}}$ instead of $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{I}}$. Unfortunately, unlike in Setting I, using resampling methods for model evaluation in Setting II results in data leakage: Although in each resampling iteration, $\mathcal{D}_{\text{test}}$ is not involved in training $\hat{f}^{\mathcal{D}^{\prime}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}}$ (the model trained on $\mathcal{D}^{\prime}_{\text{train}}$ for evaluation purposes), it is used in the tuning process performed on $\mathcal{D}_{\text{train}}$ (including $\mathcal{D}_{\text{test}}$) to obtain $\boldsymbol{\lambda}^{\mathrm{II}}$. Accordingly, since not every model generation step resulting in $\hat{f}^{\mathcal{D}^{\prime}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}}$ is conducted exclusively on $\mathcal{D}^{\prime}_{\text{train}}$, information from $\mathcal{D}_{\text{test}}$ is available during the model generation process (specifically, during tuning). Based on the definition given in Section 2.4.2, this constitutes a form of data leakage and may result in an optimistically biased resampling error (Wainer and Cawley,, 2021; Hosseini et al.,, 2020). While the inadequacy of the apparent error is widely recognized, the described pitfall associated with the resampling error is less well known and will go undetected by those not involved in model development if HP tuning is not reported (Hosseini et al.,, 2020; Lones,, 2024).
@@
As stated in Section 4.1.3, the extent to which overtuning occurs depends on the specific tuning procedure. If the HP selection is mildly overtuned, the prediction error estimate obtained from the tuning process may only exhibit a slight optimistic bias. However, as an extreme case, we can again consider the experiments from Section 4.1.3 in which HP tuning has been performed on null data (Hosseini et al.,, 2020; Varma and Simon,, 2006; Boulesteix and Strobl,, 2009; Bischl et al.,, 2023). Here, the difference between the prediction error estimate of the selected HP configuration and the true prediction error indicating random prediction is substantial, and adopting the former as the final evaluation result for a useless prediction model is clearly a biased approach.
@@
The optimistic bias of the resampling error arises because, in each resampling iteration, not all steps of the model generation process are performed exclusively on $\mathcal{D}^{\prime}_{\text{train}}$. A natural extension, therefore, is to ensure that the complete model generation is applied only to $\mathcal{D}^{\prime}_{\text{train}}$ in every iteration (see Figure 3, model evaluation c). Specifically, this implies that the tuning process is not only performed once on $\mathcal{D}_{\text{train}}$ in order to generate the final prediction model, but also on every $\mathcal{D}^{\prime}_{\text{train}}$ specified during resampling (for evaluation purposes). If the tuning process itself is based on a resampling method (i.e. if tuning is not performed using the apparent error, which is hardly ever the case if the currently described model evaluation procedure is employed), this results in two nested resampling methods. Accordingly, this procedure is called nested resampling, where the resampling method that initially splits $\mathcal{D}_{\text{train}}$ into $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ is the outer resampling loop and the resampling method creating additional splits within each $\mathcal{D}^{\prime}_{\text{train}}$ (resulting in subsets denoted as $\mathcal{D}^{\prime\prime}_{\text{train}}$ and $\mathcal{D}^{\prime}_{\text{test}}$) is the inner resampling loop (e.g., Bischl et al.,, 2023; Hosseini et al.,, 2020; Wainer and Cawley,, 2021). To distinguish nested resampling from the resampling methods discussed above and in Section 3.2.1, we will refer to the latter as simple resampling where necessary.
@@
The most straightforward form of nested resampling is the nested holdout method, where $\mathcal{D}_{\text{train}}$ is split once into $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$, and $\mathcal{D}^{\prime}_{\text{train}}$ is further divided into $\mathcal{D}^{\prime\prime}_{\text{train}}$ and $\mathcal{D}^{\prime}_{\text{test}}$. In this setup, the best HP configuration for $\mathcal{D}^{\prime}_{\text{train}}$ is determined by training and evaluating a model for each candidate HP configuration on $\mathcal{D}^{\prime\prime}_{\text{train}}$ (for training) and $\mathcal{D}^{\prime}_{\text{test}}$ (for prediction error estimation). We denote this configuration as $\boldsymbol{\lambda}^{\prime\mathrm{II}}$, as it may differ from the final prediction model’s configuration, $\boldsymbol{\lambda}^{\mathrm{II}}$, which has been obtained by tuning the model on $\mathcal{D}_{\text{train}}$ rather than $\mathcal{D}^{\prime}_{\text{train}}$. Using the HP configuration $\boldsymbol{\lambda}^{\prime\mathrm{II}}$, the model is then trained on $\mathcal{D}^{\prime}_{\text{train}}$ and evaluated on $\mathcal{D}_{\text{test}}$, which has remained unseen throughout the entire model generation process. Note that nested holdout is commonly referred to as train-validation-test split (Bischl et al.,, 2023), which, using the notation above, could also be referred to as $\mathcal{D}^{\prime\prime}_{\text{train}}$-$\mathcal{D}^{\prime}_{\text{test}}$-$\mathcal{D}_{\text{test}}$-split. Instead of holdout, any other resampling method can be used for inner and outer resampling, and it is also possible to combine different resampling methods. For example, $k$-fold CV can be used for outer resampling and holdout for inner resampling, since in the inner resampling, precise prediction error estimation is less critical as long as a sufficiently good $\boldsymbol{\lambda}^{\prime\mathrm{II}}$ is selected in each iteration (Bischl et al.,, 2023; Hosseini et al.,, 2020).
@@
While nested resampling prevents data leakage, it also has several disadvantages. First, it can be very computationally expensive, since the tuning process, which can already be time-consuming when conducted once, has to be repeated for each $\mathcal{D}^{\prime}_{\text{train}}$ specified by the outer resampling loop (Wainer and Cawley,, 2021; Bischl et al.,, 2023). Second, it is usually not feasible to conduct nested resampling with manual tuning. Apart from being even more time-demanding than nested resampling with automated tuning, it is often not possible to repeat the same tuning procedure more than once due to the informal nature of manual tuning (e.g., the user might not remember which candidate HP configurations have been evaluated during tuning). Third, like simple resampling, nested resampling does not provide an estimate of the prediction error for the final model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}}$. However, while both methods evaluate models trained on $\mathcal{D}^{\prime}_{\text{train}}$ rather than $\mathcal{D}_{\text{train}}$ (with $n^{\prime}_{\text{train}}<n_{\text{train}}$), simple resampling at least uses the same HP configuration $\boldsymbol{\lambda}^{\mathrm{II}}$ as the final prediction model. In contrast, nested resampling does not necessarily evaluate models with the same HP configuration, as each inner resampling loop may select a different configuration (see the nested holdout example above, which evaluates a model based on $\boldsymbol{\lambda}^{\prime\mathrm{II}}$ instead of $\boldsymbol{\lambda}^{\mathrm{II}}$). This makes the nested resampling result more difficult to interpret (Hosseini et al.,, 2020). The described disadvantages could explain why nested resampling estimates are not commonly reported in studies presenting new prediction models, as indicated by a recent systematic review on clinical prediction models (Andaur Navarro et al.,, 2023).

### v2

@@ 4.1.3 Tuning procedures
When tuning an HP $\lambda_{j}$, it is often not reasonable to consider all possible HP values (i.e. all values in $\Lambda_{j}$). For example, this applies if certain values of $\lambda_{j}$ are already known to cause overfitting or convergence issues. Moreover, when $\lambda_{j}$ is a preprocessing HP, $\Lambda_{j}$ may not even be formally specified (see Section 2.3.3). To perform HP tuning, it is thus essential to specify a search space $\tilde{\Lambda}_{j}$ for each HP, where $\tilde{\Lambda}_{j}$ is a bounded subset of $\Lambda_{j}$ and determines the HP values that are considered for tuning [78]. For example, if the HPs of the CART algorithm, $\lambda_{cp}$ and $\lambda_{minsplit}$ with $\Lambda_{cp}=[0,1]$ and $\Lambda_{minbucket}=\{1,\dots,n_{\text{train}}\}$, are tuned, their search spaces could be defined as $\tilde{\Lambda}_{cp}=[0.001,0.1]$ and $\tilde{\Lambda}_{minbucket}=\{5,\dots,25\}$. The (overall) search space of all $J$ HPs is denoted as $\boldsymbol{\tilde{\Lambda}}=\tilde{\Lambda}_{1}\times\dots\times\tilde{\Lambda}_{J}$.
@@
Choosing a search space involves the following trade-off: If the search space is too small, the prediction error achieved by $\boldsymbol{\tilde{\lambda}}^{\ast}$ and $\boldsymbol{\lambda}^{\ast}$ may differ greatly. On the other hand, if the search space is too large, this decreases the chance of finding $\boldsymbol{\tilde{\lambda}}^{\ast}$ (or a HP configuration that leads to a comparable prediction error) within a given computational budget [78].
@@
Unless the specified search space $\boldsymbol{\tilde{\Lambda}}$ is very small, such as when only a few categorical HPs are tuned, evaluating all HP configurations in the search space can be computationally challenging or even infeasible. For example, even if $\lambda_{cp}$ and $\lambda_{minbucket}$ are the only HPs being tuned, with the search spaces as specified above and $\tilde{\Lambda}_{cp}$ being searched in increments of 0.001, $C=100\times 21=2{,}100$ candidate HP configurations would need to be evaluated. Accordingly, one or several criteria must be specified to terminate the tuning process once it is met. The trade-off to consider when choosing a termination criterion is that the tuning process should neither stop before finding $\boldsymbol{\tilde{\lambda}}^{\ast}$ nor should it continue longer than necessary, which would result in an inefficient use of resources and, as we will discuss below, increase the risk of overtuning [78].
In automated tuning procedures, commonly used criteria are based on the number of evaluations or the runtime. However, additional criteria such as reaching a certain performance level or stagnation of performance might also be reasonable [76, 78]. Similar termination criteria, though often more intuitive than formally specified, may also exist for manual tuning when, for example, the user stops searching when satisfied by the reached performance level or gives up searching after a certain amount of time.
@@
Since, in many cases, only a subset of all HP configurations in the search space can be evaluated before the tuning process is terminated, the way in which the sequence of evaluations is determined, also called search strategy or HPO algorithm [94, 78], is another important component of the tuning procedure. Search strategies can be characterized by several aspects, such as the amount of time they spend inferring new candidate HP configurations from already evaluated ones (known as the inference vs. search trade-off; [78]). For example, search strategies such as evolutionary algorithms and Bayesian optimization consider the distribution and results of previously evaluated HP configurations to propose new configurations. In contrast, the commonly used random search strategy simply draws HP configurations from a predefined, typically uniform, distribution without taking into account past evaluations (see, e.g., [95], [78], or [76], for more details and other search strategies). In the special case where only the set of selected features is tuned, a well-known automated search strategy is backward or forward feature selection (see, e.g., [98]).
@@
In automated tuning procedures, all HPs are usually tuned jointly, i.e. each evaluated HP configuration potentially considers different values of each HP. However, the HPs could also be tuned sequentially, i.e. the complete tuning procedure is repeated for each HP [124, 140]. For example, in a setting with three HPs (i.e. $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\lambda_{3})$), $\lambda_{1}$ would be tuned first with $\lambda_{2}$ and $\lambda_{3}$ set to default, which yields $\lambda_{1}^{\mathrm{II}}$. Then, $\lambda_{2}$ is tuned with $\lambda_{1}=\lambda_{1}^{\mathrm{II}}$ and $\lambda_{3}$ set to its default. Finally, $\lambda_{3}$ is tuned with $\lambda_{1}=\lambda_{1}^{\mathrm{II}}$ and $\lambda_{2}=\lambda_{2}^{\mathrm{II}}$, yielding $\lambda_{3}^{\mathrm{II}}$. As sequential tuning does not consider any interaction effects between the HPs, it is generally less likely to yield a $\boldsymbol{\lambda}^{\mathrm{II}}$ comparable to $\boldsymbol{\tilde{\lambda}}^{\ast}$ than joint tuning. On the other hand, sequential tuning demands less time, with the maximum number of evaluations increasing linearly rather than exponentially with the number of HPs to tune, as is the case with joint tuning. Hence, it could be a realistic approach for manual tuning.
@@
Due to the optimistic bias of the apparent error, the standard approach for automated HP tuning is to employ a resampling method. In the case of $k$-fold CV, which is a common choice for HP tuning [78], this means that for each candidate HP configuration $\boldsymbol{\lambda}^{(c)}$, $k$ models are trained and evaluated on different subsets of $\mathcal{D}_{\text{train}}$.
@@
While resampling methods provide an improvement over using the apparent error, the corresponding estimators also exhibit a certain degree of pessimistic bias and variance (with the degree of bias and variance depending on the resampling method used, as discussed in Section 3.2.1). A potential pitfall arising from the variance is that the winning HP configuration, $\boldsymbol{\lambda}^{\mathrm{II}}$, may have been selected simply because the trained prediction model(s) using $\boldsymbol{\lambda}^{\mathrm{II}}$ performed particularly well by chance on the specified test data set(s) $\mathcal{D}_{\text{test}}$, which are the same for each evaluated HP configuration. This means that the HP selection has essentially been overfitted to the respective test data set(s) $\mathcal{D}_{\text{test}}$, which in this context is also referred to as overtuning, overhyping, or oversearching [104, 125, 120, 95, 78, 83]. If the true prediction error of $\boldsymbol{\lambda}^{\mathrm{II}}$ is still comparable to the prediction error of $\boldsymbol{\tilde{\lambda}}^{\ast}$, overtuning effects are negligible. However, there might also be scenarios in which the true prediction error of $\boldsymbol{\lambda}^{\mathrm{II}}$ is no better, or even worse, than that of the default HP configuration, but its estimated prediction error is drastically deflated (i.e. over-optimistic), as the corresponding prediction model(s) that were trained during resampling incidentally fit very well to the specific noise pattern in the respective test data set(s) $\mathcal{D}_{\text{test}}$. This has been demonstrated in several experiments where tuning was conducted on null data (i.e. data without any true signal), yet the prediction error estimate of the selected HP configuration $\boldsymbol{\lambda}^{\mathrm{II}}$ was substantially smaller (i.e. better) than its true prediction error indicating random prediction [104, 138, 80, 78].
@@
Note that since the HPs are overfitted to the test data set(s) $\mathcal{D}_{\text{test}}$, which are not seen during training on the corresponding $\mathcal{D}^{\prime}_{\text{train}}$, overtuning occurs on a higher level than overfitting of the model parameters (see Section 3.1). Accordingly, overtuning effects may only be visible after evaluating a large number of HP configurations [78]. However, literature suggests that the risk of overtuning does not only depend on the number of evaluated HP configurations but also, for example, on the search strategy, the type of tuned HP, and the number of observations in $\mathcal{D}_{\text{train}}$ [104, 139, 83]. In general, overtuning is considered an open problem of HP tuning, and although strategies have been suggested to avoid it (e.g., using different splits for each evaluation, [118]), there are no commonly agreed-upon solutions [95].
@@ 4.2.1 Evaluation of a model generated on all available data
As in Setting I, reporting the apparent error for model evaluation is inappropriate in Setting II (see Figure 3, model evaluation a). In this case, however, the designated test data set $\mathcal{D}_{\text{test}}=\mathcal{D}_{\text{train}}=\mathcal{D}$ is even used twice during model generation: first during the HP tuning process and then again during the final training process. Depending on the specific tuning procedure employed, this can introduce an even greater optimistic bias compared to, for example, using default HP values. Although the apparent error is generally not suitable for assessing a model’s performance, some users who performed tuning via resampling may mistakenly believe it now reflects a form of resampling error. This was noted by [119], who also reference a paper that appears to have fallen into this pitfall.
@@
Similar to Setting I, an alternative evaluation procedure in Setting II is to employ a resampling method (see Figure 3, model evaluation b). In principle, the chosen resampling method is carried out as described in Section 3.2.1, except that in each resampling iteration, the model is trained on $\mathcal{D}^{\prime}_{\text{train}}$ and evaluated on $\mathcal{D}_{\text{test}}$ with $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{II}}$ instead of $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{I}}$. Unfortunately, unlike in Setting I, using resampling methods for model evaluation in Setting II results in data leakage: Although in each resampling iteration, $\mathcal{D}_{\text{test}}$ is not involved in training $\hat{f}^{\mathcal{D}^{\prime}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}}$ (the model trained on $\mathcal{D}^{\prime}_{\text{train}}$ for evaluation purposes), it is used in the tuning process performed on $\mathcal{D}_{\text{train}}$ (including $\mathcal{D}_{\text{test}}$) to obtain $\boldsymbol{\lambda}^{\mathrm{II}}$. Accordingly, since not every model generation step resulting in $\hat{f}^{\mathcal{D}^{\prime}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}}$ is conducted exclusively on $\mathcal{D}^{\prime}_{\text{train}}$, information from $\mathcal{D}_{\text{test}}$ is available during the model generation process (specifically, during tuning). Based on the definition given in Section 2.4.2, this constitutes a form of data leakage and may result in an optimistically biased resampling error [139, 104]. While the inadequacy of the apparent error is widely recognized, the described pitfall associated with the resampling error is less well known and will go undetected by those not involved in model development if HP tuning is not reported [104, 114].
@@
As stated in Section 4.1.3, the extent to which overtuning occurs depends on the specific tuning procedure. If the HP selection is mildly overtuned, the prediction error estimate obtained from the tuning process may only exhibit a slight optimistic bias. However, as an extreme case, we can again consider the experiments from Section 4.1.3 in which HP tuning has been performed on null data [104, 138, 80, 78]. Here, the difference between the prediction error estimate of the selected HP configuration and the true prediction error indicating random prediction is substantial, and adopting the former as the final evaluation result for a useless prediction model is clearly a biased approach.
@@
The optimistic bias of the resampling error arises because, in each resampling iteration, not all steps of the model generation process are performed exclusively on $\mathcal{D}^{\prime}_{\text{train}}$. A natural extension, therefore, is to ensure that the complete model generation is applied only to $\mathcal{D}^{\prime}_{\text{train}}$ in every iteration (see Figure 3, model evaluation c). Specifically, this implies that the tuning process is not only performed once on $\mathcal{D}_{\text{train}}$ in order to generate the final prediction model but also on every $\mathcal{D}^{\prime}_{\text{train}}$ specified during resampling (for evaluation purposes). If the tuning process itself is based on a resampling method (i.e. if tuning is not performed using the apparent error, which is hardly ever the case if the currently described model evaluation procedure is employed), this results in two nested resampling methods. Accordingly, this procedure is called nested resampling, where the resampling method that initially splits $\mathcal{D}_{\text{train}}$ into $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ is the outer resampling loop and the resampling method creating additional splits within each $\mathcal{D}^{\prime}_{\text{train}}$ (resulting in subsets denoted as $\mathcal{D}^{\prime\prime}_{\text{train}}$ and $\mathcal{D}^{\prime}_{\text{test}}$) is the inner resampling loop [78, 104, 139, e.g.,]. To distinguish nested resampling from the resampling methods discussed above and in Section 3.2.1, we will refer to the latter as simple resampling where necessary.
@@
The most straightforward form of nested resampling is the nested holdout method, where $\mathcal{D}_{\text{train}}$ is split once into $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$, and $\mathcal{D}^{\prime}_{\text{train}}$ is further divided into $\mathcal{D}^{\prime\prime}_{\text{train}}$ and $\mathcal{D}^{\prime}_{\text{test}}$. In this setup, the best HP configuration for $\mathcal{D}^{\prime}_{\text{train}}$ is determined by training and evaluating a model for each candidate HP configuration on $\mathcal{D}^{\prime\prime}_{\text{train}}$ (for training) and $\mathcal{D}^{\prime}_{\text{test}}$ (for prediction error estimation). We denote this configuration as $\boldsymbol{\lambda}^{\prime\mathrm{II}}$, as it may differ from the final prediction model’s configuration, $\boldsymbol{\lambda}^{\mathrm{II}}$, which has been obtained by tuning the model on $\mathcal{D}_{\text{train}}$ rather than $\mathcal{D}^{\prime}_{\text{train}}$. Using the HP configuration $\boldsymbol{\lambda}^{\prime\mathrm{II}}$, the model is then trained on $\mathcal{D}^{\prime}_{\text{train}}$ and evaluated on $\mathcal{D}_{\text{test}}$, which has remained unseen throughout the entire model generation process. Note that nested holdout is commonly referred to as train-validation-test split [78], which, using the notation above, could also be referred to as $\mathcal{D}^{\prime\prime}_{\text{train}}$-$\mathcal{D}^{\prime}_{\text{test}}$-$\mathcal{D}_{\text{test}}$-split. Instead of holdout, any other resampling method can be used for inner and outer resampling, and it is also possible to combine different resampling methods. For example, $k$-fold CV can be used for outer resampling and holdout for inner resampling, since in the inner resampling, precise prediction error estimation is less critical as long as a sufficiently good $\boldsymbol{\lambda}^{\prime\mathrm{II}}$ is selected in each iteration [78, 104].
@@
While nested resampling prevents data leakage, it also has several disadvantages. First, it can be very computationally expensive, since the tuning process, which can already be time-consuming when conducted once, has to be repeated for each $\mathcal{D}^{\prime}_{\text{train}}$ specified by the outer resampling loop [139, 78]. Second, it is usually not feasible to conduct nested resampling with manual tuning. Apart from being even more time-demanding than nested resampling with automated tuning, it is often not possible to repeat the same tuning procedure more than once due to the informal nature of manual tuning (e.g., the user might not remember which candidate HP configurations have been evaluated during tuning). Third, like simple resampling, nested resampling does not provide an estimate of the prediction error for the final model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}}$. However, while both methods evaluate models trained on $\mathcal{D}^{\prime}_{\text{train}}$ rather than $\mathcal{D}_{\text{train}}$ (with $n^{\prime}_{\text{train}}<n_{\text{train}}$), simple resampling at least uses the same HP configuration $\boldsymbol{\lambda}^{\mathrm{II}}$ as the final prediction model. In contrast, nested resampling does not necessarily evaluate models with the same HP configuration, as each inner resampling loop may select a different configuration (see the nested holdout example above, which evaluates a model based on $\boldsymbol{\lambda}^{\prime\mathrm{II}}$ instead of $\boldsymbol{\lambda}^{\mathrm{II}}$). This makes the nested resampling result more difficult to interpret [104]. The described disadvantages could explain why nested resampling estimates are not commonly reported in studies presenting new prediction models, as indicated by a recent systematic review on clinical prediction models [74].